OpenStudy (anonymous):

The Parthenon temple of ancient Greece has a rectangular shape. It measures almost 70 meters long and 30 meters wide. How would the area of the temple change if its length is halved?

OpenStudy (mathsolver):

the area would also be halved

OpenStudy (anonymous):

about an hour agoLuRockzz StHilaire A It's area reduces to half of the original C Its area reduces to one fourth of the original D Its area reduces to one sixteenth of the original B It's area reduces to the square root of the original

OpenStudy (anonymous):

Why would it be halved?

OpenStudy (mathsolver):

let the width and length be a and b (it doesnt matter which way round) and the area of the temple be A. since a*b=A then (a/2)*b=A/2

OpenStudy (anonymous):

Ohhh, okay! Thanks!